# NLP-ADBench: NLP Anomaly Detection Benchmark

### v1

@@
Abstract.
Anomaly detection (AD) is a critical machine learning task with diverse applications in web systems, including fraud detection, content moderation, and user behavior analysis. Despite its significance, AD in natural language processing (NLP) remains underexplored, limiting advancements in detecting anomalies in text data such as harmful content, phishing attempts, or spam reviews. In this paper, we introduce NLP-ADBench, the most comprehensive benchmark for NLP anomaly detection (NLP-AD), comprising eight curated datasets and evaluations of nineteen state-of-the-art algorithms. These include three end-to-end methods and sixteen two-step algorithms that apply traditional anomaly detection techniques to language embeddings generated by bert-base-uncased and OpenAI’s text-embedding-3-large models.
Our results reveal critical insights and future directions for NLP-AD. Notably, no single model excels across all datasets, highlighting the need for automated model selection. Moreover, two-step methods leveraging transformer-based embeddings consistently outperform specialized end-to-end approaches, with OpenAI embeddings demonstrating superior performance over BERT embeddings. By releasing NLP-ADBench at https://github.com/USC-FORTIS/NLP-ADBench, we provide a standardized framework for evaluating NLP-AD methods, fostering the development of innovative approaches. Th is work fills a crucial gap in the field and establishes a foundation for advancing NLP anomaly detection, particularly in the context of improving the safety and reliability of web-based systems.
Anomaly Detection, Natural Language Processing, Web Mining, NLP Benchmark, Benchmark, Datasets
*Equal contribution.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Anomaly detectionccs: Information systems Web miningccs: Computing methodologies Natural language processing
1. Introduction
@@
Anomaly detection (AD) is a fundamental area in machine learning with diverse applications in web systems, such as fraud detection, content moderation, and user behavior analysis (Chandola et al., 2009; Ahmed et al., 2016). While significant progress has been made in AD for structured data types, including tabular, graph, and time series data (Chalapathy and Chawla, 2019; Han et al., 2022; Lai et al., 2021; Liu et al., 2022), its application to natural language processing (NLP) tasks—a domain crucial to many web-based applications—remains both essential and relatively under explored (Ruff et al., 2021). For instance, detecting harmful or abusive language is critical for ensuring safe user experiences in social media platforms and online forums (Fortuna and Nunes, 2018). Similarly, in e-commerce, identifying anomalies in reviews or product descriptions can help detect spam or fraudulent behaviors, critical for preserving user trust and platform integrity.
Key Contributions. In this work, we introduce NLP-ADBench, a comprehensive benchmark for NLP anomaly detection (NLP-AD). Building on our prior benchmarking efforts in tabular (Han et al., 2022), graph (Liu et al., 2022), and time-series (Lai et al., 2021) anomaly detection, NLP-ADBench (as the most comprehensive NLP-AD benchmark) provides several key advantages over existing works (Bejan et al., 2023): (i) a diverse collection of 8 datasets reflecting real-world web scenarios, (ii) integration of 19 advanced NLP-based anomaly detection methods, (iii) actionable insights for advancing NLP-AD research, and (iv) fully open-sourced datasets, algorithm implementation, and experiments.
@@
Key Insights/Takeaways (see details in §3). Our experiments reveal the following key findings: (i) There is no single model that consistently performs best across all datasets due to variability in dataset characteristics, highlighting the need for automated model selection. (ii) Transformer-based embeddings significantly enhance the performance of two-step AD methods, such as LUNAR (Goodge et al., 2022) and LOF (Breunig et al., 2000), compared to end-to-end approaches. (iii) High-dimensional embeddings, such as those from OpenAI models, improve detection accuracy but require balancing performance with computational efficiency. (iv) Addressing dataset-specific biases and reducing human dependency in anomaly definitions remain critical for advancing robust and generalizable NLP-AD systems.
Our study aligns closely with the Web Mining and Content Analysis” track, particularly addressing topics such as Robustness and generalizability of Web mining methods” and “Web data quality in the era of algorithmically-generated content.”
2. NLP-ADBench: AD Benchmark for NLP Tasks
2.1. Preliminaries and Problem Definition
@@
Anomaly Detection in Natural Language Processing (NLP-AD) focuses on identifying text instances that deviate significantly from expected or typical patterns. Unlike structured data, text data is inherently unstructured, high-dimensional, and deeply influenced by the nuances of human language, including syntax, semantics, and context (Aggarwal, 2017).These unique properties introduce significant challenges, making the development of robust and accurate AD methods for NLP a complex and demanding task.
@@
Formally, let $\mathcal{D}={x_{1},x_{2},\dots,x_{N}}$ denote a corpus where each $x_{i}$ is a text instance. The goal of NLP-AD is to learn an anomaly scoring function $f:\mathcal{X}\rightarrow\mathbb{R}$ that assigns a real-valued anomaly score to each text instance. Higher scores correspond to greater deviations from normal patterns, indicating a higher likelihood of an instance being anomalous.
Table 1. Statistical information of the NLPAD dataset.
@@
NLPAD-BBCNews $1,785$ $1,723$ $62$ $3.47\%$
NLPAD-EmailSpam $3,578$ $3,432$ $146$ $4.08\%$
@@
2.2. Curated Benchmark Datasets
Each transformed dataset is named by adding the prefix “NLPAD-” to the original dataset’s name (e.g., NLPAD-AGNews, NLPAD-BBCNews), distinguishing them from the original datasets. The NLPAD datasets are provided in a unified JSON Lines format [^1] A format for storing structured data, see https://jsonlines.org/. for compatibility and ease of use. Each line is a JSON object [^2] A lightweight data format, see https://www.json.org/json-en.html. with four fields: text (the text used for anomaly detection), label (the anomaly detection label, where 1 represents an anomaly and 0 represents normal), original_task (the task of the original dataset), and original_label (the category label from the original dataset).
To transform each dataset for NLP-AD, we established a text selection process based on the data format. For tabular data, we carefully chose appropriate columns as the text source. For document-based data, we extracted text directly from relevant documents. The anomalous class for each dataset was selected based on semantic distinctions within the dataset categories, ensuring that the identified anomalies represent meaningful deviations from the normal data distribution (Emmott et al., 2015; Han et al., 2022). Once identified, the anomalous class was downsampled to represent less than 10% of the total instances.
2.3. The Most Comprehensive NLP-AD Algorithms with Open Implementations
Table 2. Performance comparison of 19 Algorithms on 8 NLPAD datasets using AUROC, with best results highlighted in bold and shaded.
@@ 2.3. The Most Comprehensive NLP-AD Algorithms with Open Implementations
Compared to the existing NLP-AD benchmark by Bejan et al. (Bejan et al., 2023), NLP-ADBench provides a broader evaluation by including 19 algorithms, categorized into two groups. The first group comprises 3 end-to-end algorithms that directly process raw text data to produce anomaly detection outcomes. The second group consists of 16 algorithms derived by applying 8 traditional anomaly detection (AD) methods to text embeddings generated from two models: bert-base-uncased (Devlin et al., 2019) and OpenAI’s text-embedding-3-large (OpenAI, 2024). These traditional AD methods do not operate on raw text directly but instead perform anomaly detection on embeddings, offering a complementary approach to the end-to-end methods. This comprehensive algorithm collection enables a robust evaluation of direct and embedding-based NLP anomaly detection techniques. Here, we provide a brief description; see details in Appx. A.2.
End-to-end NLP-AD Algorithms. We evaluate three end-to-end algorithms tailored for NLP anomaly detection. (1) Context Vector Data Description (CVDD) (Ruff et al., 2019) leverages context vectors and pre-trained embeddings with a multi-head self-attention mechanism to project normal instances close to learned contexts, identifying anomalies based on deviations. (2) Detecting Anomalies in Text via Self-Supervision of Transformers (DATE) (Manolache et al., 2021) trains transformers using self-supervised tasks like replaced mask detection to capture normal text patterns and flag anomalies. (3) Few-shot Anomaly Detection in Text with Deviation Learning (FATE) (Das et al., 2023) uses a few labeled anomalies with deviation learning to distinguish anomalies from normal instances, achieving state-of-the-art performance on benchmark datasets.
Two-step NLP-AD Algorithms. We evaluate 8 two-step algorithms that rely on embeddings generated by models such as bert-base-uncased (Devlin et al., 2019) and text-embedding-3-large (OpenAI, 2024). These algorithms are designed to work with structured numerical data and cannot directly process raw textual data, requiring text transformation into numerical embeddings. (4) LOF (Breunig et al., 2000) measures local density deviations, while (5) DeepSVDD (Ruff et al., 2018) minimizes the volume of a hypersphere enclosing normal representations. (6) ECOD (Li et al., 2022) uses empirical cumulative distribution functions to estimate densities and assumes anomalies lie in distribution tails. (7) IForest (Liu et al., 2008) recursively isolates anomalies through random splits, and (8) SO_GAAL (Liu et al., 2019) generates adversarial samples to identify anomalies. Reconstruction-based approaches include (9) AE (Aggarwal, 2017), which flags anomalies based on reconstruction errors, and (10) VAE (Kingma and Welling, 2013; Burgess et al., 2018), which identifies anomalies using reconstruction probabilities or latent deviations. Finally, (11) LUNAR (Goodge et al., 2022) enhances traditional local outlier detection with graph neural networks.
Accessibility. We open-source NLP-ADBench with datasets and algorithm implementation for reproducibility and future research.
3. Experiment Results
@@
Figure 1. Average rank on AUCROC of 19 NLPAD methods across 8 datasets (the lower the better).
3.1. Experiment Setting
Hyperparameter Settings. For all the algorithms in NLP-ADBench, we use their default hyperparameter (HP) settings in the original paper for a fair comparison, same as ADBench (Han et al., 2022).
Evaluation Metrics and Statistical Tests. We evaluate different NLP-AD methods by a widely used metric: AUCROC (Area Under Receiver Operating Characteristic Curve).
Embeddings Definitions:
  1. (1)
    BERT refers specifically to the bert-base-uncased model (Devlin et al., 2019).
  2. (2)
    OpenAI refers to OpenAI’s text-embedding-3-large model (OpenAI, 2024).
  3. (3)
3.2. Results, Discussions, and New Directions
@@
No single model performs best across all datasets due to variability in dataset characteristics. The performance of AD models varies significantly depending on dataset characteristics, making it challenging to identify a universal winner. For datasets with a larger variety of categories, such as NLPAD-AGNews, OpenAI + LUNAR achieved a score of 0.9226, outperforming the end-to-end method CVDD (0.6046) by 52.6%. Similarly, on NLPAD-BBCNews, OpenAI + LOF scored 0.9558, exceeding CVDD (0.7221) by 32.4%. Conversely, for datasets with fewer categories or binary-class structures, end-to-end methods show distinct advantages. For example, on the binary-class NLPAD-SMSSpam dataset, the end-to-end method DATE achieved 0.9398, outperforming OpenAI + LUNAR (0.7189) by 30.7%.
@@
  • Future Direction 1: Automated Model Selection. These results emphasize the importance of developing automated approaches to select the most suitable model based on the characteristics of the dataset, e.g., the number of categories and the complexity of the structure. One feasible solution will be adapting the meta-learning framework from tabular AD settings (Zhao et al., 2021) to NLPAD.
@@
Transformer-based embeddings boost the performance of two-step AD methods. Two-step AD algorithms, when paired with transformer-based embeddings, have shown the ability to outperform end-to-end methods like CVDD and FATE in NLP AD tasks. For instance, OpenAI + LUNAR achieved a significantly higher performance of 0.9226 on the NLPAD-AGNews dataset, outperforming CVDD by 52.6% and FATE by 19.0%. Similarly, OpenAI + LOF achieved 0.9558 on the NLPAD-BBCNews dataset, surpassing CVDD by 32.4% and FATE by 2.7%. These results emphasize the substantial advantage of leveraging transformer-based embeddings over relying solely on end-to-end methods. The success of two-step approaches can be attributed to the high-quality contextual embeddings produced by transformer models like OpenAI, which serve as a robust foundation for AD. In contrast, some end-to-end methods, such as CVDD, rely on embeddings from older models like GloVe, highlighting the need for incorporating more advanced embeddings to improve performance.
  • Future Direction 2: Efficient and Robust Embedding Utilization. While transformer-based embeddings deliver superior accuracy, their computational cost and resource requirements remain significant. Future research should focus on designing lightweight, efficient algorithms that can harness updated embeddings without excessive time or resource consumption. Additionally, improving the robustness of two-step approaches to ensure consistent performance across diverse datasets and anomaly types will be crucial for advancing NLP anomaly detection systems.
@@
High-dimensional embeddings improve detection but require balancing performance and efficiency. The dimensionality of embeddings varies significantly across models, impacting both performance and efficiency in AD. While the BERT-base model generates 768-dimensional embeddings, OpenAI’s text-embedding-3-large produces embeddings with a dimensionality of 3072—a 300% increase. Our experiments reveal that OpenAI embeddings consistently outperform BERT embeddings across multiple datasets in NLP-ADBench, demonstrating the advantages of higher-dimensional embeddings. For instance, OpenAI + LUNAR achieved a performance of 0.9452 on the NLPAD-YelpReview dataset, surpassing BERT + LUNAR’s 0.6522 by 44.9%. Similarly, on NLPAD-AGNews, OpenAI + LUNAR scored 0.9226, exceeding BERT + LUNAR’s 0.7694 by 19.9%. On NLPAD-N24News, OpenAI + LUNAR reached 0.8320, outperforming BERT + LUNAR’s 0.6284 by 32.4%. These results highlight the capability of high-dimensional embeddings to enhance AD performance. However, the increased dimensionality also introduces challenges, such as higher computational costs and potential information redundancy.
@@
  • Future Direction 3: Optimizing Embedding Dimensionality. Future research should explore dimensionality reduction techniques tailored to NLP-AD tasks to mitigate redundancy and computational overhead without compromising performance. Additionally, hybrid approaches that adaptively adjust dimensionality based on dataset characteristics could enable scalable, efficient, and effective AD systems.
4. Conclusion, Limitations, and Future Works

### v2

@@
Abstract
Anomaly detection (AD) is an important machine learning task with applications in fraud detection, content moderation, and user behavior analysis. However, AD is relatively understudied in a natural language processing (NLP) context, limiting its effectiveness in detecting harmful content, phishing attempts, and spam reviews. We introduce NLP-ADBench, the most comprehensive NLP anomaly detection (NLP-AD) benchmark to date, which includes eight curated datasets and 19 state-of-the-art algorithms. These span 3 end-to-end methods and 16 two-step approaches that adapt classical, non-AD methods to language embeddings from BERT and OpenAI. Our empirical results show that no single model dominates across all datasets, indicating a need for automated model selection. Moreover, two-step methods with transformer-based embeddings consistently outperform specialized end-to-end approaches, with OpenAI embeddings outperforming those of BERT. We release NLP-ADBench at https://github.com/USC-FORTIS/NLP-ADBench , providing a unified framework for NLP-AD and supporting future investigations.
NLP-ADBench: NLP Anomaly Detection Benchmark
Yuangang Li1,*, Jiaqi Li1,*, Zhuo Xiao1,*, Tiankai Yang1, Yi Nian1, Xiyang Hu2, Yue Zhao1 1University of Southern California  2Arizona State University {yuangang, jli77629, zhuoxiao, tiankaiy, yinian, yzhao010}@usc.edu, xiyanghu@asu.edu Corresponding authors
1 Introduction
@@
Anomaly detection (AD) is a fundamental area in machine learning with diverse applications in web systems, such as fraud detection, content moderation, and user behavior analysis Chandola et al. (2009); Ahmed et al. (2016). Substantial progress has been achieved in AD for structured data such as tabular, graph, and time series Chalapathy and Chawla (2019); Han et al. (2022); Lai et al. (2021); Liu et al. (2022), but its extension to natural language processing (NLP) remains relatively underexplored Ruff et al. (2021); Yang et al. (2024). This gap limits our ability to identify harmful content, phishing attempts, and spam reviews.
For instance, detecting abusive or threatening language is crucial for ensuring that social media platforms and online forums remain safe environments for users Fortuna and Nunes (2018). Likewise, detecting anomalous product reviews or descriptions in e-commerce is important for preserving user trust and platform credibility Chino et al. (2017). However, many standard AD methods are designed for numeric or categorical data and are not easily adapted to unstructured text Zhao et al. (2019); Chen et al. (2024). Existing studies on NLP-specific AD are limited in both dataset variety and algorithmic range Han et al. (2022); Liu et al. (2022); Yang et al. (2024), leaving open questions about which approaches work best under different conditions. These gaps lead to a central research question: How can we systematically evaluate and compare diverse AD methods across real-world text datasets, and what insights can be gained to guide future development in NLP-based AD?
Our Proposal and Key Contributions. We introduce NLP-ADBench, the most comprehensive benchmark for NLP-AD tasks. NLP-ADBench offers four major benefits compared to prior work Bejan et al. (2023): (i) eight real-world datasets covering a wide range of web use cases; (ii) 19 advanced methods that apply standard AD algorithms to language embeddings or use end-to-end neural architectures; (iii) detailed empirical findings that highlight new directions for NLP-AD; and (iv) fully open-source resources, including datasets, algorithm implementations, and more, aligns with the Resources and Evaluation track.
@@
Key Insights/Takeaways (see details in §3). Our comprehensive experiments reveal: (i) No single model dominates across all datasets, showing the need for model selection; (ii) Transformer-based embeddings substantially boost two-step AD methods (e.g., LUNAR Goodge et al. (2022) and LOF Breunig et al. (2000)) relative to end-to-end approaches; (iii) High-dimensional embeddings (e.g., from OpenAI) improve detection performance, but also raise computational overhead; and (iv) Dataset-specific biases and human-centered anomaly definitions remain challenging for building robust and widely applicable NLP-AD systems.
2 NLP-ADBench: AD Benchmark for NLP Tasks
2.1 Preliminaries and Problem Definition
@@
Anomaly Detection in Natural Language Processing (NLP-AD) focuses on identifying text instances that deviate significantly from expected or typical patterns. Unlike structured data, text data is inherently unstructured, high-dimensional, and deeply influenced by the nuances of human language, including syntax, semantics, and context Aggarwal (2017); Yang et al. (2024). These unique properties introduce significant challenges, making the development of robust and accurate AD methods for NLP a complex and demanding task.
@@
Formally, let $\mathcal{D}={x_{1},x_{2},\dots,x_{N}}$ denote a corpus where each $x_{i}$ is a text instance. The goal of NLP-AD is to learn an anomaly scoring function $f:\mathcal{X}\rightarrow\mathbb{R}$ that assigns a real-valued anomaly score to each text instance. Higher scores denote greater deviations from normal patterns, indicating a higher likelihood of an anomalous instance.
2.2 Curated Benchmark Datasets
Each transformed dataset is named by adding the prefix “NLPAD-" to the original dataset’s name (e.g., NLPAD-AGNews, NLPAD-BBCNews), distinguishing them from the original datasets. The NLPAD datasets are provided in a unified JSON Lines format for compatibility and ease of use. Each line is a JSON object with four fields: text (the text used for anomaly detection), label (the anomaly detection label, where 1 represents an anomaly and 0 represents normal), original_task (the task of the original dataset), and original_label (the category label from the original dataset).
To transform each dataset for NLP-AD, we established a text selection process based on the data format. For tabular data, we carefully chose appropriate columns as the text source. For document-based data, we extracted text directly from relevant documents. The anomalous class for each dataset was selected based on semantic distinctions within the dataset categories, ensuring that the identified anomalies represent meaningful deviations from the normal data distribution Emmott et al. (2015); Han et al. (2022). Once identified, the anomalous class was downsampled to represent less than 10% of the total instances.
Table 1: Statistical information of the NLPAD dataset.
@@ 2.2 Curated Benchmark Datasets
NLPAD-BBCNews $1,785$ $1,723$ 6262 $3.47\%$
NLPAD-EmailSpam $3,578$ $3,432$ 146146 $4.08\%$
@@
2.3 The Most Comprehensive NLP-AD Algorithms with Open Implementations
Compared to the existing NLP-AD benchmark by Bejan et al. Bejan et al. (2023), NLP-ADBench provides a broader evaluation by including 19 algorithms, categorized into two groups. The first group comprises 3 end-to-end algorithms that directly process raw text data to produce anomaly detection outcomes. The second group consists of 16 algorithms derived by applying 8 traditional anomaly detection (AD) methods to text embeddings generated from two models: bert-base-uncased Devlin et al. (2019) and OpenAI’s text-embedding-3-large OpenAI (2024). These traditional AD methods do not operate on raw text directly but instead perform anomaly detection on embeddings, offering a complementary approach to the end-to-end methods. This comprehensive algorithm collection enables a robust evaluation of direct and embedding-based NLP anomaly detection techniques. Here, we provide a brief description; see details in Appx. A.2.
End-to-end NLP-AD Algorithms. We evaluate 3 end-to-end algorithms tailored for NLP-AD. (1) Context Vector Data Description (CVDD) Ruff et al. (2019) leverages context vectors and pre-trained embeddings with a multi-head self-attention mechanism to project normal instances close to learned contexts, identifying anomalies based on deviations. (2) Detecting Anomalies in Text via Self-Supervision of Transformers (DATE) Manolache et al. (2021) trains transformers using self-supervised tasks like replaced mask detection to capture normal text patterns and flag anomalies. (3) Few-shot Anomaly Detection in Text with Deviation Learning (FATE) Das et al. (2023) uses a few labeled anomalies with deviation learning to distinguish anomalies from normal instances. We adapt it to train solely on normal data, referring to the adapted version as FATE*.
Two-step NLP-AD Algorithms. We evaluate 8 two-step algorithms that rely on embeddings generated by models such as bert-base-uncased Devlin et al. (2019) and text-embedding-3-large OpenAI (2024). These algorithms are designed to work with structured numerical data and cannot directly process raw textual data, requiring text transformation into numerical embeddings. (4) LOF Breunig et al. (2000) measures local density deviations, while (5) DeepSVDD Ruff et al. (2018) minimizes the volume of a hypersphere enclosing normal representations. (6) ECOD Li et al. (2022) uses empirical cumulative distribution functions to estimate densities and assumes anomalies lie in distribution tails. (7) IForest Liu et al. (2008) recursively isolates anomalies through random splits, and (8) SO_GAAL Liu et al. (2019) generates adversarial samples to identify anomalies. Reconstruction-based approaches include (9) AE Aggarwal (2017), which flags anomalies based on reconstruction errors, and (10) VAE Kingma and Welling (2013); Burgess et al. (2018), which identifies anomalies using reconstruction probabilities or latent deviations. Finally, (11) LUNAR Goodge et al. (2022) enhances traditional local outlier detection with graph neural networks.
3 Experiment Results
3.1 Experiment Setting
Hyperparameter Settings. For all the algorithms in NLP-ADBench, we use their default hyperparameter (HP) settings in the original paper for a fair comparison, same as ADBench Han et al. (2022).
Evaluation Metrics and Statistical Tests. We evaluate different NLP-AD methods by a widely used metric: AUROC (Area Under Receiver Operating Characteristic Curve) and AUPRC (Area Under Precision-Recall Curve) value.
Embeddings Definitions:
  1. 1.
    BERT refers specifically to the bert-base-uncased model Devlin et al. (2019).
  2. 2.
    OpenAI refers to OpenAI’s text-embedding-3-large model OpenAI (2024).
  3. 3.
3.2 Results, Discussions, and New Directions
We analyze the AUROC results presented in Table 2 and the average rank summary in Figure 1. For completeness, AUPRC scores and their corresponding average ranks are reported in Appendix A.3.
Table 2: Performance comparison of 19 Algorithms on 8 NLPAD datasets using AUROC, with best results highlighted in bold and shaded.
@@ 3.2 Results, Discussions, and New Directions
Figure 1: Average rank on AUROC of 19 NLPAD methods across 8 datasets (the lower the better).
@@
No single model consistently excels across all datasets due to variability in dataset characteristics. AD model performance varies significantly across datasets, complicating the selection of a universally optimal model. For datasets with more categories (e.g., NLPAD-AGNews), two-step methods like OpenAI + LUNAR (0.9226) outperform end-to-end methods such as CVDD (0.6046) by 52.6%. Similarly, on NLPAD-BBCNews, OpenAI + LOF (0.9558) surpasses CVDD (0.7221) by 32.4%. Conversely, on the binary-class datasets (e.g., NLPAD-SMSSpam), end-to-end methods perform better, with DATE (0.9398) clearly exceeding OpenAI + LUNAR (0.7189) by 30.7%.
@@
  • Future Direction 1: Automated Model Selection. These results emphasize the importance of developing automated approaches to select the most suitable model. One feasible solution will be adapting the meta-learning framework from tabular AD settings Zhao et al. (2021) to NLP-AD.
@@
Transformer-based embeddings boost the performance of two-step AD methods. Two-step AD algorithms paired with transformer-based embeddings have consistently outperformed end-to-end methods in NLP-AD tasks. For instance, OpenAI + LUNAR achieves 0.9226 on NLPAD-AGNews, surpassing CVDD by 52.6% and FATE* by 19.0%. Similarly, OpenAI + LOF reaches 0.9558 on NLPAD-BBCNews, exceeding CVDD by 32.4% and FATE* by 2.7%. This advantage arises primarily because two-step methods leverage superior contextual embeddings from modern transformer models (e.g., OpenAI), whereas end-to-end methods like CVDD rely on older embeddings (e.g., GloVe). This highlights the need for end-to-end methods to adopt more advanced embeddings to enhance performance.
  • Future Direction 2: Transformer Embedding Integration for End-to-End AD. Future end-to-end methods should adopt transformer-based embeddings over static embeddings like GloVe. Research should focus on embedding integration optimized for end-to-end AD frameworks.
@@
High-dimensional embeddings enhance detection but require balancing performance and efficiency. Embedding dimensionality significantly impacts both performance and computational efficiency in AD tasks. Compared to BERT-base embeddings (768 dimensions), OpenAI’s text-embedding-3-large embeddings (3072 dimensions, a 300% increase) consistently achieve superior results across multiple datasets in NLP-ADBench. Specifically, OpenAI + LUNAR achieves 0.9452 on NLPAD-YelpReview (outperforming BERT + LUNAR’s 0.6522 by 44.9%), 0.9226 on NLPAD-AGNews (exceeding BERT + LUNAR’s 0.7694 by 19.9%), and 0.8320 on NLPAD-N24News (surpassing BERT + LUNAR’s 0.6284 by 32.4%). These results clearly demonstrate the advantage of higher-dimensional embeddings for enhancing AD performance. However, higher dimensionality also introduces greater computational costs and potential information redundancy.
@@
  • Future Direction 3: Optimizing Embedding Dimensionality. Future research should explore NLP-AD-specific dimensionality reduction techniques to reduce redundancy and computational costs without compromising performance. Additionally, adaptive methods that dynamically adjust dimensionality based on dataset characteristics could enhance scalability and efficiency.
3.3 In-depth Analysis of Key Findings
3.3.1 Explaining Method-Dataset Fit
To understand why certain models outperform others on specific datasets, we conduct both quantitative and qualitative analyses to identify dataset-level factors influencing model performance.
Quantitative Corpus-Level Linguistic Analysis We characterize each dataset using three linguistic indicators: (1) Avg-Len, the average number of BERT tokens per sample (text complexity); (2) Lexical-Burstiness, the proportion of top-20 anomaly-specific tokens from BERT-tokenized TF-IDF (higher values indicate stronger lexical anomaly signals); (3) Topic-Diversity, the Shannon entropy over the label distribution of normal-class examples (0 for binary datasets; higher values indicate broader topic coverage).
Table 3: Dataset characteristics and best method
Dataset Avg-Len Lexical-Burst. Topic-Div. Best Method
NLPAD-AGNews 39.5 0.040 1.585 OpenAI+LUNAR (two-step)
NLPAD-BBCNews 481.7 0.028 1.981 OpenAI+LUNAR (two-step)
NLPAD-EmailSpam 238.4 0.024 0.000 DATE (end-to-end)
NLPAD-Emotion 20.3 0.087 1.949 OpenAI+LUNAR (two-step)
NLPAD-MovieReview 290.3 0.026 0.000 OpenAI+ECOD (two-step)
NLPAD-N24News 1034.2 0.016 4.437 OpenAI+LUNAR (two-step)
NLPAD-SMSSpam 20.5 0.063 0.000 DATE (end-to-end)
NLPAD-YelpReview 151.6 0.023 0.000 OpenAI+LUNAR (two-step)
From Table. 3, we get two findings: (1) Datasets with short texts, high lexical-burstiness, and explicit lexical markers (e.g., NLPAD-SMSSpam) tend to favor DATE’s token-level anomaly detection. (2) Datasets with moderate length, lower lexical-burstiness, and high topical diversity (e.g., NLPAD-AGNews) benefit from richer, context-sensitive embeddings (OpenAI+LUNAR).
Qualitative Analysis of Representative Cases To further investigate the model performance gap, we qualitatively examined anomaly examples from NLPAD-SMSSpam and NLPAD-AGNews (see Table 4). In NLPAD-SMSSpam, anomalies often contain explicit lexical irregularities, such as numeric tokens, unconventional formatting, or urgency-inducing phrases. These surface-level features align closely with DATE’s self-supervised scoring mechanism, which is sensitive to token-level deviations. By contrast, anomalies in NLPAD-AGNews exhibit subtle semantic shifts without distinctive lexical markers. Detecting such anomalies requires a deeper understanding of contextual semantics, which exceeds DATE’s capacity and favors two-step methods with embedding-based models such as OpenAI and LUNAR.
Table 4: Representative anomaly examples from NLPAD-SMSSpam and NLPAD-AGNews
NLPAD-SMSSpam NLPAD-AGNews
PRIVATE! Your 2003 Account Statement for shows 800 un redeemed S. I. M. points. Call 08715203694 Identifier Code: 40533 Expires 31/10/04 NEW YORK U.S. stocks are expected to open modestly higher Tuesday as investors use a slight let up in the rise in oil prices to add to portfolios…
FREE for 1st week! No1 Nokia tone 4 ur mob every week just txt NOKIA to 87077 Get txting and tell ur mates. zed POBox 36504 W45WQ norm150p/tone 16
Urgent! Please call 09061213237 from landline. 5000 cash or a luxury 4 Canary Islands Holiday await collection. T Cs SAE PO Box 177. M227XY. 150ppm. 16 AFP Indian shares, Asia’s second top performers last year, are poised for long term gains as foreign investors buy into the market, seeing the country as an economic "growth story," according to analysts.
XXXMobileMovieClub: To use your credit, click the WAP link in the next txt message or click here xxxmobilemovieclub.com?n QJKGIGHJJGCBL
3.3.2 Performance-Efficiency Trade-offs in Embedding Dimensionality
While OpenAI-based two-step methods achieve high anomaly detection performance, their high dimensionality raises concerns about computational and financial cost in deployment scenarios. To explore whether such overhead is justified, we conduct dimensionality reduction experiments using PCA, projecting OpenAI embeddings to 768 dimensions—the same as BERT.
Table 5: Performance-efficiency trade-offs of dimensionality reduction on datasets
Method
Embedding Dim
Performance Runtime (s/sample)
AUROC AUPRC Total Embedding PCA Inference
NLPAD-AGNews
OpenAI + LUNAR (orig.) 3072 0.907 0.562 0.377 0.267 0.000 0.110
OpenAI + LUNAR (PCA) 768 0.890 0.546 0.409 0.252 0.007 0.150
BERT + LUNAR 768 0.790 0.325 0.081 0.056 0.000 0.025
OpenAI + LOF (orig.) 3072 0.896 0.575 0.314 0.239 0.000 0.075
OpenAI + LOF (PCA) 768 0.798 0.321 0.435 0.245 0.007 0.183
BERT + LOF 768 0.771 0.304 0.076 0.062 0.000 0.014
NLPAD-MovieReview
OpenAI + LUNAR (orig.) 3072 0.664 0.238 0.322 0.270 0.000 0.052
OpenAI + LUNAR (PCA) 768 0.681 0.249 0.409 0.270 0.007 0.132
BERT + LUNAR 768 0.467 0.152 0.071 0.060 0.000 0.010
OpenAI + LOF (orig.) 3072 0.652 0.242 0.314 0.274 0.000 0.041
OpenAI + LOF (PCA) 768 0.624 0.226 0.384 0.256 0.007 0.121
BERT + LOF 768 0.498 0.166 0.068 0.060 0.000 0.009
As shown in Table 5, PCA-reduced OpenAI embeddings slightly change performance (e.g., AUROC drops from 0.907 to 0.890 on NLPAD-AGNews with LUNAR) and consistently outperform BERT at the same dimension. However, PCA increases total runtime because it compresses embeddings into a denser space, which complicates decision boundaries and slows down inference.
This finding reinforces Future Direction 3, highlighting the need for NLP-AD-specific dimensionality reduction techniques that balance representation quality with computational efficiency.
4 Conclusion